OpenStudy (anonymous):

what is radical 6r times radical 3r

hero (hero):

\[\sqrt{3r} \times \sqrt{6r} = \sqrt{(6r)(3r)}\]

hero (hero):

Finish simplifying from there @memy

hero (hero):

By the way, In general, \[\sqrt{a} \times \sqrt{b} = \sqrt{ab}\]

OpenStudy (anonymous):

\[\sqrt{18r ^{2}} = \sqrt{18 \times r ^{2}} = \sqrt{9\times2\times r ^{2}} \]

OpenStudy (anonymous):

I made a mistake sorry :)

OpenStudy (anonymous):

and can you do it now?

OpenStudy (anonymous):

thank you hero and pfeh. 1999 for all your help. i understand now

OpenStudy (anonymous):

welcome ;)

hero (hero):

@memy do you know what it simplifies to? @PFEH.1999 gave you a really good hint.

OpenStudy (anonymous):

i was thinking 3r squared radical 2 ut i may be wrong

hero (hero):

Doesn't square and square root cancel? By the way radicals are the same as square root.

hero (hero):

For example \[\sqrt{a^2} = \sqrt[2]{a^2} = a\] The square and the root cancels leaving just a.

hero (hero):

So if you have \[\sqrt{r^2}\] Do you know what that simplifies to?

OpenStudy (anonymous):

no im confused now sorry not so good in math lol

hero (hero):

As a said earlier, a radical is the same as a square root. Square roots and squares are inverses of each other. Two things that are inverses of each other usually cancel out.

OpenStudy (anonymous):

last thing i got was 9 times 2 times r sqaured ll under the radical so i know the 9 turns into the three but the two isnt a perfect square so doesnt that have to stay under the radical/

hero (hero):

Yes, but now we need to figure out what to do with the \(r^2\) that's under the radical

OpenStudy (anonymous):

do we leave ring it out with the three and leave the two alone under the radical

hero (hero):

You reduced \(\sqrt{9}\) to 3 because \(\sqrt{9}\) = \(\sqrt{3^2} = 3\)

hero (hero):

Again, the root and the square also canceled in that case.

hero (hero):

\(\sqrt{3^2} = 3\) \(\sqrt{r^2} = ?\)

OpenStudy (anonymous):

ok so if the radicals and square roots cancel out are we left with just 3 radical two oh ok now i understand so would it e 3r radical 2 then

OpenStudy (anonymous):

and it's good for you to know: \[\sqrt{a ^{2}} = \pm a\]

OpenStudy (anonymous):

r

OpenStudy (anonymous):

for example: \[\sqrt{9} = \pm3 \]

hero (hero):

It does, but @PFEH.1999, that doesn't apply in this case. We're only concerned with positive numbers since r implies radius and the expression implies distance.

OpenStudy (anonymous):

yes,yes you're quite write! \[\sqrt{9} = \pm3 \] but: \[\sqrt{r ^{2}} = | \pm r|\]

OpenStudy (anonymous):

sorry right not write

hero (hero):

Either way, @memy has it correct now.

OpenStudy (anonymous):

thnk you oth for your help tonight really appreciate it